David is X years old and Martin is 3 years older if the sum of their ages is 37 write an equation based on this information

Answer by jim_thompson5910(35256)

David's current age = x
Martin's current age = x+3

The sum of their ages has us add up the two expressions (x and x+3) so we end up with (x)+(x+3) = x+x+3 = 2x+3

We're told that "the sum of their ages is 37", so set 2x+3 equal to 37 to get the answer 2x+3 = 37